OpenStudy (anonymous):

which equation describes a line that is NOT parallel to the lines described by the other three equations? 1. 2x+7y=8 2. -6x-21y=-5 3.-10.5y=3x 4.y+8=(2/7)(x-7)

OpenStudy (anonymous):

OH MY GAWD

OpenStudy (anonymous):

the struggle!

OpenStudy (mathstudent55):

Can you check the third equation?

OpenStudy (anonymous):

Yup thats the exact one on my test. its -10.5y=3x

OpenStudy (mathstudent55):

Ok. Solve each equation for y and write it in the form y = mx + b. m, the number that multiplies x, is the slope.

OpenStudy (mathstudent55):

Let's try the first one: \(2x+7y=8\) Subtract 2x from both sides: \(7y = -2x + 8\) Divide both sides by 7: \(y = -\dfrac{2}{7} x + \dfrac{8}{7} \) The slope is \(-\dfrac{2}{7} \) Now do the same to all other equations and compare their slopes.

OpenStudy (mathstudent55):

By "do the same to all other equations" I mean solve all other equations for y and put them in the y = mx + b form.

OpenStudy (anonymous):

ok so 2x+7y=8 is y=-2x/7+8/7 -6x-21y=-5 is y=-2x/7+5/21 -10.5y=3x is y=-0.28571429x and y+8=2/7(x-7) is y=2x/7-10

OpenStudy (anonymous):

1. slope is -2/7 2. slope is -2/7 3. slope is -0.28571429 4. slope is 2/7

OpenStudy (anonymous):

so is it 3? @mathstudent55

OpenStudy (mathstudent55):

\(-6x-21y=-5\) \(-21y = 6x - 5\) \(y = -\dfrac{6}{21}x + \dfrac{5}{21} \) \(y = -\dfrac{2}{7}x + \dfrac{5}{21} \) Correct.

OpenStudy (anonymous):

Thankyou so much for your help!

OpenStudy (mathstudent55):

\(-10.5y=3x\) \(y = - \dfrac{3}{10.5} x \) \(y = - \dfrac{6}{21} x \) \(y = - \dfrac{2}{7} x \) Incorrect.

OpenStudy (mathstudent55):

You should keep the third one as a fraction. The you will see that the third one is also -2/7. The answer is D, 2/7

OpenStudy (mathstudent55):

1. slope is -2/7 2. slope is -2/7 3. slope is -0.28571429 = -2/7 4. slope is 2/7 The only different slope is D.